If you bought $4000 worth of some items and sold it for $5000 dollars, making $50 on each, how much the each Item were originally bought
Answer by arobin4(1)

let 'a' be the quantity of items purchased.
let 'b' be the original cost of each item.
(a)(b) = 4000 This is Equation 1
(a)(b + 50) = 5000 This is Equation 2
Solve Equation 1 and Equation 2 simultaneously.
Step 1: Solve for 'a' in Equation 1 to get (a)=4000/(b)
Step 2: Substitute 'a' from step 1 for 'a' in step 2 to get Equation 3 below
(4000 / b)(b + 50) = 5000 This is Equation 3
Step 3: Solve Equation 3 for 'b' to get b=$200
'b' is the original cost of each item. Therefore the answer is $200.